### v1

# CookAR: Affordance Augmentations in Wearable AR to Support Kitchen Tool Interactions for People with Low Vision

Abstract.
Cooking is a central activity of daily living, supporting independence and both mental and physical health. However, prior work has highlighted key barriers for people with low vision (LV) to cook, particularly around safely interacting with cooking tools, such as sharp knives or hot pans. Drawing on recent advancements in computer vision (CV), we present CookAR, a head-mounted AR system with real-time object affordance augmentations to support safe and efficient interactions with kitchen tools. To design and implement CookAR, we manually collected and annotated the first egocentric dataset of kitchen tool affordances, fine-tuned an affordance segmentation model, and developed an AR system with a stereo camera to generate visual augmentations. To validate CookAR, we conducted a technical evaluation of our fine-tuned model as well as a qualitative lab study with 10 LV participants for suitable augmentation design. Our technical evaluation demonstrates that our model outperforms the baseline on our tool affordance dataset, while our user study indicates a preference for affordance augmentations over the traditional whole object augmentations. Code is available at: https://github.com/makeabilitylab/CookAR.
1. Introduction
Cooking is an essential activity of daily living, supporting independence (Miyawaki et al., 2012; Mynatt and Rogers, 2001; Taylor et al., 2016) and both mental and physical health (Mills et al., 2017; Bilyk et al., 2009; Jones et al., 2019; Taylor et al., 2016). However, cooking also involves significant visual tasks that can be challenging or dangerous for blind and low vision (BLV) people, especially when interacting with kitchen tools, such as sharp knives or hot pans (Bilyk et al., 2009; Jones et al., 2019; Li et al., 2021, 2024; Wang et al., 2023b).
Unlike those who are completely blind, people with low vision (LV)—vision loss that cannot be corrected using glasses or contact lenses (Corn and Erin, 2010)—often rely on their residual vision in daily activities and use different low vision tools to enhance visual information (Szpiro et al., 2016a, b). With recent advancements in AI-powered augmented reality (AR) technology, researchers have explored new possibilities for supporting LV individuals by automatically recognizing their environment and providing appropriate visual augmentations. Prior AR research prototypes have been developed to assist with tasks such as visual search (Zhao et al., 2016), stair navigation (Zhao et al., 2019), and sports (Lee et al., 2023). However, none have been specifically designed to support meal preparation. Moreover, prior AR systems mainly focus on understanding effective augmentation designs (Zhao et al., 2019), often oversimplifying the computer vision (CV) recognition in their development, thus neglecting the effects of technological limitations (e.g., CV inaccuracies and system delays) on user experience.
We introduce CookAR, a fully-functional wearable stereo AR prototype that recognizes and augments the affordances of cooking tools in real-time to support LV individuals with efficient and safe interactions in the kitchen. In contrast to prior research that augments objects as a whole (Zhao et al., 2016; Fox et al., 2023), we distinguish and augment the object affordance (i.e., component parts that afford interactions), such as the safe-to-handle “grabbable” areas and the dangerous-to-touch “hazardous” areas (Figure 1). To enable accurate affordance recognition, we constructed a custom egocentric image dataset for kitchen tool affordances by selecting and labeling images from the Epic Kitchens dataset (Damen et al., 2018) and fine-tuned an affordance segmentation model. We then leveraged stereo depth estimation using the ZED Mini111https://store.stereolabs.com/products/zed-mini stereo camera and an Oculus Quest 2222https://www.meta.com/quest/products/quest-2/ headset with video passthrough to precisely overlay affordance augmentations on the 3D environment in real-time.
To evaluate CookAR, we conducted a technical evaluation of our fine-tuned model as well as a three-part qualitative lab study with 10 LV participants to evaluate their experiences with CookAR and solicit feedback about potential affordance augmentation designs. For the model assessment, we found that our fine-tuned affordance segmentation model (mAP of 46.3%) outperformed the base RTMDet (Lyu et al., 2022) model (mAP of 12.3%) in accurate tool affordance recognition and segmentation. For the three-part user study, LV participants were first asked to locate and pickup cooking tools across three conditions: (1) without CookAR (baseline), (2) with CookAR displaying whole object augmentations, and (3) with CookAR displaying affordance augmentations. They then completed a free-form cooking task with CookAR (Part 2) and brainstormed desired augmentation designs using design probes (Part 3). Findings indicate that participants prefer affordance augmentations over whole object augmentations in a kitchen, as they enable faster understanding of an object’s spatial arrangement and safe interaction parts. Most participants preferred affordance augmentations consisting of green solid overlay on grabbable areas and red outlines on hazardous areas. Moreover, participants identified five additional tool affordances with desired augmentations, including entry (e.g., cup rim), exit (e.g., carafe spout), containment (e.g., cup base), intersection (e.g., knife blade on butter), and activation (e.g., carafe buttons) areas, all of which should be outlined in a contrasting color (e.g., black or white).
In summary, our research contributions include: (1) CookAR, a fully-functional AI-powered wearable AR prototype that augments kitchen tool affordances for low vision users to enable safe and efficient tool interactions; (2) an egocentric affordance dataset for kitchen tools and an accompanying fine-tuned affordance segmentation model; and (3) user study results with 10 LV participants that reveal user experiences with CookAR, preferences for augmentation designs, and five newly desired affordance areas.
2. Related Work
Our work builds on prior formative research exploring low-vision cooking, wearable AR to enhance accessibility, and affordance segmentation.
2.1. Challenges in Low Vision Cooking
Low vision people face various challenges in activities of daily living, such as cooking (Bilyk et al., 2009; Wang et al., 2023a, b; Li et al., 2021, 2024), shopping (Szpiro et al., 2016b), navigation (Zhao et al., 2018; Szpiro et al., 2016b), and sports (Lee et al., 2023; Rector et al., 2015). Among these activities, cooking is an essential task for an independent and healthy life (Taylor et al., 2016). However, cooking also poses major barriers and safety concerns to blind and low vision (BLV) people as they need to interact with various ingredients and kitchen tools, such as sharp knives and hot pans (Kashyap, 2020; Kim et al., 2022). As a result, BLV people tend to eat more pre-processed food or frequently dine at restaurants, which can negatively impact their health (Jones et al., 2019; Montero, 2005).
To better understand how BLV people engage in cooking tasks, prior work has conducted both interview and observational studies (Li et al., 2021; Lin et al., 2014; Wang et al., 2023a, b). For example, Jones et al. (Jones et al., 2019) surveyed 101 BLV participants in the U.K. about their shopping and cooking experiences, revealing that vision loss made cooking difficult and the difficulty level was correlated with the severity of visual impairments. Li et al. (Li et al., 2021) analyzed 122 YouTube videos of BLV people preparing meals and interviewed 12 BLV participants about their cooking experiences. They identified several cooking-related challenges, such as utilizing cooking tools and tracking object dynamics in the kitchen. A follow-up contextual inquiry study (Li et al., 2024) examined how BLV people recognize cookware and utensils and measure ingredients. The study identified essential cooking-related information to convey, such as position, safety, and orientation information.
Specifically for low-vision (LV) people, Wang et al. (Wang et al., 2023a) conducted a contextual inquiry study, observing and comparing the cooking experiences between six LV participants and four blind participants. They found that while blind participants relied on touch, LV people extensively used their vision during cooking. However, compared to blind people, LV people felt less confident, less safe, and more tired and stressed during cooking. Moreover, they were less satisfied with the currently-available cooking tools than blind people, indicating a need for technology that considers low vision people’s unique needs. The study also identified key challenges faced by LV people, such as distinguishing objects with low contrast and safely interacting with dangerous kitchen tools. Wang et al. (Wang et al., 2023b) further interviewed six low vision rehabilitation professionals to understand current training strategies and tools to support cooking. However, they highlighted that current solutions cannot fully address all cooking challenges faced by LV people. Our research fills this gap by building an AR system to support safe and efficient cooking tool interactions for low vision people by providing visual augmentations.
2.2. Using Wearable AR to Enhance Accessibility
In the fields of accessibility and HCI, wearable AR has been used to support people with diverse disabilities, for example, AR systems that caption and visualize speech and sounds for deaf or hard of hearing (DHH) individuals (Findlater et al., 2019; Jain et al., 2018a, b, 2015; Miller et al., 2017; Olwal et al., 2020; Peng et al., 2018; Schipper and Brinkman, 2017), support hands-free interactions with screen displays for people with upper body motor impairments (McNaney et al., 2014; Malu and Findlater, 2014, 2015), offer speech support for people with aphasia (Williams et al., 2015), and provide social cue therapy for children with autism (Washington et al., 2017).
Within the low-vision aid context, head-worn AR devices can selectively enhance user’s vision by interpreting their environment and tasks (Azenkot and Zhao, 2017; Zhao et al., 2017). For instance, prior research has developed AR systems that can capture real-time video feeds of the surroundings and apply image processing techniques (Massof and Rickman, 1992; Zhao et al., 2016; Deemer et al., 2018) to enhance visual information, such as edge enhancement (Massof et al., 1994; Kwon et al., 2012; Hwang and Peli, 2014), scene recoloring based on distance (Everingham et al., 1999; Hicks et al., 2013; van Rheede et al., 2015), and pixel remapping for visual field loss (Loshin and Juday, 1989; Gupta et al., 2018; Peli, 2001; Luo and Peli, 2006). However, while these solutions are beneficial for simple tasks like reading (Stearns et al., 2017; Huang et al., 2019; Deemer et al., 2018), they still lack semantic understanding of the scene and cannot effectively support more complex activities involving object interactions, such as cooking (Wang et al., 2023b).
More recently, researchers have combined AR and CV to develop scene-aware visual augmentations aimed at assisting LV people with more intricate activities like visual search (Zhao et al., 2016), stair navigation (Zhao et al., 2019), wayfinding (Zhao et al., 2020), obstacle maneuver (Fox et al., 2023), button pressing (Lang and Machulla, 2021), and sports (Lee et al., 2023). Nonetheless, no prior systems have addressed the unique cooking challenge that involves dynamic tool interactions. Moreover, prior AR research for low vision mainly focuses on designing and evaluating suitable visual augmentations for low vision people. They tend to oversimplify the CV recognition in system development, for example, using QR codes (Zhao et al., 2016, 2019) or existing spatial mapping APIs (Zhao et al., 2020; Fox et al., 2023) to anchor augmentations to the real-world environment. As a result, they overlook the technical challenges in building a real-time AI-powered AR system, as well as the potential impact of the technological limitations (e.g., CV errors, system latency) on users’ experiences. Our research addresses this gap by developing CookAR, a fully-functional wearable AR system that recognizes and augments kitchen tool affordances in real-time to support safe and efficient interactions in a cooking scenario.
2.3. Affordance Segmentation
In contrast to most prior research that augments whole objects (Zhao et al., 2016; Fox et al., 2023; Lee et al., 2023), our research focuses on recognizing and augmenting tool affordances. Affordance is traditionally defined as “the opportunities for actions that objects offer, relative to the user’s ability to perceive and act on them” (Gibson, 1966, 1977, 2014). Highlighting object affordances can effectively guide human attention and actions (Garrido-Vásquez and Schubö, 2014; Tünnermann et al., 2015; Rehrig et al., 2022). Despite the prominence of affordance segmentation in robotics (Chu et al., 2019a, b; Nguyen et al., 2016, 2017; Bahl et al., 2023) and computer vision (Luddecke and Worgotter, 2017; Chuang et al., 2018; Luo et al., 2022), automatic affordance augmentation has received comparatively less attention and applications in the field of HCI. Notably, there exists a gap for an affordance dataset in the context of a kitchen environment, especially designed for the needs of individuals with low vision. To address this, we built an AR system that focuses on affordance segmentation and enhancement. We created a new dataset focused on the affordances of kitchen tools by selecting and annotating image frames from the egocentric Epic Kitchens dataset (Damen et al., 2018) and fine-tuned an instance segmentation model on our dataset.
3. System Implementation
We designed and built CookAR, an innovative solution that integrates AR and CV technologies to improve object interactions for individuals with low vision. Unlike traditional enhancements that target objects as a whole, our prototype highlights their affordances (i.e., functional parts), facilitating easier identification and interaction with areas to grasp or avoid. To create a fully-functional wearable AR system, we needed to address both the computer vision problem of accurately recognizing object affordances in real-time and the HCI problem of designing and rendering suitable affordance augmentations. In this section, we describe our approach for each step, including (1) collecting and annotating a dataset focused on the affordances of kitchen tools; (2) fine-tuning an instance segmentation and recognition model on our dataset to detect these affordances; and (3) building a head-mounted AR system with a stereo camera to render visual augmentations based on the outputs of our affordance recognition model.
3.1. Data Collection and Annotation
To train an effective AI model for affordance segmentation, we first need an appropriate dataset. However, to the best of our knowledge, there is no prior cooking tool dataset with annotations to enable affordance segmentation and recognition. Below, we describe our multi-step process to collect and annotate object affordances in egocentric cooking images. This labeled dataset is one contribution of our work and will be open-sourced upon paper acceptance to enable future research.
Data Collection. To build a kitchen tool affordance image dataset, we used an egocentric video dataset called Epic Kitchens (Damen et al., 2018), which consists of 100 hours of video footage of sighted people cooking in their homes. We selected the Epic Kitchens video dataset since it not only involves a wide range of cooking scenarios with various kitchen tools, but also captures video feeds from the first-person perspective, which aligns with the egocentric nature of head-worn AR devices.
Because the Epic dataset is so large, we needed to filter critical frames of interest. We used YOLOv8 (Jocher et al., 2023) trained on the MS COCO dataset (Lin et al., 2014) to detect and collect frames featuring cooking-related objects, such as spoons, knives, forks, cups, scissors, sinks, and dining tables. We skipped 20 frames after finding at least one of those objects to minimize repetition. We then manually reviewed the selected frames to empirically remove similar, excessively blurry, or irrelevant images, resulting in 4,928 key frames.
Data Annotation & Augmentation. We then labeled these frames using the image annotation feature in Roboflow333https://roboflow.com, a platform that facilitates building and deploying end-to-end computer vision models by offering tools for annotation, training, and optimization. Roboflow enables labeling automation using the Segment Anything model (SAM) (Kirillov et al., 2023), allowing us to easily select and segment the interactive parts of objects (e.g., knife blade vs. knife handle) and add corresponding class labels.
Drawing on prior work in low-vision cooking (Li et al., 2021, 2024; Wang et al., 2023b), the research team collectively decided on 18 distinct classes: Knife Blade, Knife Handle, Spoon Bowl, Spoon Handle, Fork Tines, Fork Handle, Scissor Blade, Scissor Handle, Ladle Bowl, Ladle Handle, Spatula Head, Spatula Handle, Pan Base, Pan Handle, Cup Base, Cup Handle, Carafe Base, and Carafe Handle (Figure 2). When labeling, we adhered to the following heuristic: (1) the object should visually resemble the class it is being labeled as; and (2) the object should serve functions similar to those of the label class. For instance, a large wooden spoon can be tricky to label, as it can resemble a spoon, ladle, or spatula, and have versatile use such as stirring a pan (like a spoon), scooping contents from a pot (like a ladle), or lifting eggs (like a spatula) across different images in the dataset. We labeled these ambiguous objects based on their shape and use in a given frame. The annotations were cross-checked between six members of the research team to reduce errors and bias.
After annotating, we used various image augmentation techniques available on Roboflow to enhance the dataset for better generalizability across real-world scenarios, including: cropping with 0% minimum zoom and 40% maximum zoom, rotation between -15° and +15°, brightness between -15% and +15%, blur up to 2.5px, and noise up to 0.1% of pixels. We then adjusted the images to fit a 640x480 resolution (i.e., the MS COCO average image resolution (Lin et al., 2014)) to accommodate our chosen base model’s preferences and facilitate their use in future research. This resulted in a final dataset of 10,152 images. The entire dataset is publicly available here: https://github.com/makeabilitylab/CookAR.
3.2. Model Fine-Tuning
To provide real-time object affordance information to LV users, we fine-tuned the RTMDet model (Lyu et al., 2022), specifically its RTMDet-Ins-l variant, on our kitchen tool affordance dataset. This model is the current state-of-the-art in real-time instance segmentation444https://paperswithcode.com/sota/real-time-instance-segmentation-on-mscoco, offering robust accuracy and 300+ FPS on an NVIDIA 3090 GPU. RTMDet features large kernel depthwise convolution and batch normalization layers, pre-trained on MS COCO.
We opted to fine-tune RTMDet instead of training it from scratch, which allowed us to better leverage our smaller, class-specific dataset. To achieve this, we initialized the base RTMDet-Ins-l model with pre-trained weights, froze its backbone, and adjusted the model configuration file for our label classes, before training it on our kitchen tool affordance dataset. This customized model, dubbed RTMDet-Ins-l-Cook, was trained over 150 epochs with a batch size of 4 on a single CUDA-enabled NVIDIA 4080 GPU. Because RTMDet-Ins-l-Cook underwent fine-tuning on a dataset with affordance annotations, it can mimic an affordance segmentation model’s capabilities while retaining RTMDet’s real-time performance. We provide a technical evaluation in Section 4.
3.3. The CookAR Prototype
With our fine-tuned RTMDet-Ins-l-Cook model, we built CookAR, a wearable AR prototype that can recognize and visually augment the affordances of kitchen tools in real-time. To implement CookAR, we needed to address key technical and HCI challenges, including: (1) how to spatially highlight object affordances in 3D space; (2) how to run a real-time model capable of providing affordance feedback with minimal latency; (3) and how to best visually indicate affordances to LV users to support but not overwhelm their existing perceptual vision systems.
To generate visual augmentations that align with the object parts in a 3D space, we built a custom stereo video see-through AR system by combining the ZED Mini stereo camera with an Oculus Quest 2 VR headset, as off-the-shelf AR headsets (e.g., Microsoft HoloLens 2) do not yet support long range real-time depth sensing and heavy CV models. The affordance visualizations themselves are rendered as colored polygon overlays. Specifically, we used green (hexcode #3BE8B0) to indicate a graspable area and red (#FC626A) to indicate a risky area. See Figure 1. As our user study section describes (Section 5), we also further explored and brainstormed other affordance augmentation designs.
Because our real-time CV model is computationally expensive, CookAR is tethered to a laptop with a NVIDIA 4080 mobile GPU. The CookAR system first captures image frames using the ZED Mini stereo camera and streams them to an external server via the Transmission Control Protocol (TCP) for affordance segmentation by our RTMDet-Ins-l-Cook model (confidence threshold of 0.4). Then, it converts the resulting JSON with affordance masks and labels into a Protocol Buffers message555https://protobuf.dev for efficient streaming. This message is then sent back to be processed by the ZED Mini API (Stereolabs, 2024), which deserializes the message back into a JSON and creates a ZED-compatible texture (or colored overlay) for each affordance mask. Finally, the ZED Mini performs stereo depth estimation and overlays the textures onto the left and right image frames for binocular vision in the Quest 2 headset. While the ZED Mini stereo camera allowed us to prototype CookAR, it required our system to be tethered to a computer, which we alleviated by using lengthy cables.
In our latency analysis, we ran CookAR for five minutes and computed the average latency of each component: video streaming from ZED to computer took 16.76ms; affordance recognition took 15.95ms; result streaming back to ZED took 10.43ms; and depth estimation and augmentation rendering took 3.39ms. All other components had negligible impact on runtime. The overall latency is on average 46.82ms per frame (21.36 FPS), resulting in a near real-time system. System structure is summarized in Figure 3.
4. Technical Evaluation
We first conducted a technical evaluation of our fine-tuned RTMDet-Ins-l-Cook model, comparing its performance against the base RTMDet-Ins-l model on our kitchen tool affordance dataset. Findings indicate that our fine-tuned model is significantly more accurate in recognizing and segmenting affordances of cooking than the unmodified model.
4.1. Methods
To assess the performance of the base and fine-tuned models on our kitchen tool affordance dataset, we leveraged the model testing pipeline provided by the MMDetection library (Chen et al., 2019), a PyTorch-based open-source toolkit for object detection and segmentation, which performs evaluations using the test subset of a given dataset. Using Roboflow, we generated a testing set of 596 images with an 82-12-6 train-validation-test split and ensured that our fine-tuned model was not exposed to images in the test subset.
For instance segmentation tasks, accuracy is conventionally measured using three key metrics: segmentation mean average precision (mAP), AP at a 50% Intersection over Union (IOU) threshold (AP@50), and AP at a 75% IOU threshold (AP@75) (He et al., 2017). Each metric serves a distinct purpose:
mAP offers a comprehensive assessment of a model’s precision across various detection thresholds, by averaging precision at multiple recall levels for each class. It also aggregates results across a range of Intersection over Union (IoU) thresholds, from 0.5 to 0.95 in steps of 0.05, providing a holistic view of model performance across different degrees of overlap between the predicted masks and the ground truth;
AP@50 focuses on the precision of segmentation at a specific IoU threshold of 50%, a more lenient measure that considers a prediction correct if the overlap with the ground truth is at least half;
AP@75 evaluates precision at a stricter IoU threshold of 75%, demanding higher accuracy in the overlap between the predicted segmentation and the ground truth for a positive assessment.
IoU, central to these metrics, quantitatively evaluates the overlap between predicted segmentation masks and the actual ground truth, serving as a direct measure of accuracy in spatial alignment. We applied these metrics to compare the performance of our fine-tuned model against the baseline, aiming to capture the nuances of improvement across different levels of strictness in segmentation accuracy. We computed and compared the three metrics across the base and fine-tuned models.
| Model Name | mAP | AP@50 | AP@75 |
| RTMDet-Ins-l (on COCO) | 0.437 | 0.660 | 0.470 |
| RTMDet-Ins-l (on our dataset) | 0.123 | 0.199 | 0.310 |
| RTMDet-Ins-l-Cook (on our dataset) | 0.463 | 0.749 | 0.486 |
4.2. Results
Our findings (Table 1) show that the base RTMDet-Ins-l model performs poorly on our affordance-specific dataset despite its competency on the COCO dataset (Lin et al., 2014) of 43.7% segmentation mAP. Conversely, our fine-tuned RTMDet-Ins-l-Cook model excelled in identifying and segmenting cooking tool components, demonstrating a significantly higher segmentation mAP of 46.3%, compared to the base model’s 12.3%. This improvement was also evident in our model’s performance at different IoU thresholds, with AP@50 and AP@75 reaching 74.9% and 48.6%, significantly outperforming the base model’s 19.9% and 13.2%. These results highlight the enhanced accuracy of our fine-tuned RTMDet-Ins-l-Cook model on our kitchen tool affordance dataset.
In addition, we show several inference results of RTMDet-Ins-l-Cook on images from the test subset of our dataset in Figure 4. Our model demonstrates impressive robustness, identifying and segmenting graspable, safe areas even when hands or other partial occlusions are present in the images.
5. User Study
We then evaluated our CookAR prototype in a three-part qualitative lab study with 10 low vision participants. As initial work, our goals were to evaluate how LV participants might benefit from a real-time object affordance augmentations to complete cooking tasks, solicit their reactions to a fully-functional but early-stage prototype (e.g., how do they react to augmentation errors), and co-brainstorm visual overlay designs via design probes. Participants provided feedback throughout the study and answered open-ended questions regarding their experiences, which were recorded and transcribed for later analysis.
5.1. Participants
To achieve a diverse participant pool, we recruited 10 LV participants from two different cities via mailing lists and snowball sampling. Participants were screened using a demographic questionnaire, which gathered information on age, gender, vision condition, and prior experience with AR and AI technologies. The average age was 62.2 years (), with a gender distribution of 70% female and 30% male. Participants had a broad range of low vision conditions with visual acuity ranging from 20/40 to 20/400 and visual field loss at different areas—see Table 2. Most participants reported little to no experience with AR and AI, except for P1 who had used both technologies.
| P# | Gender | Age | Left Eye Acuity | Right Eye Acuity | Description of Visual Field |
| P1 | Male | 30 | No Light Perception | 20/400 | Coloboma dominates the right superior portion of my right eye. |
| P2 | Female | 83 | 20/200 | 20/100 | Deteriorating eyesight from dry macular degeneration. Lost central vision on left eye. Central vision on the right eye is still there but not good. Have peripheral vision on both. |
| P3 | Female | 62 | 20/125 | 20/100 | Low vision. Some holes in it, like black spots. Scar tissue. |
| P4 | Male | 65 | 20/20 | 20/60 | Can see from 2/3’s of left eye, some far right peripheral vision from right eye. |
| P5 | Female | 70 | 20/200 | 20/100 | Macular degeneration and side effects of chemotherapy. Blurry vision and need font enlargement to read. Visual field intact. |
| P6 | Male | 50 | 20/40 | No Light Perception | Blind in right eye. Need glasses for left. Visual field intact. |
| P7 | Female | 81 | 20/200 | 20/60 | Diminished vision due to macular degeneration. Visual field intact. |
| P8 | Female | 80 | 20/50 | 20/50 | Have dry macular degeneration with loss of some vision in the center of my left eye. |
| P9 | Female | 30 | Light Perception | 20/80 | Can make out faces with right eye. Left eye is blind. Visual field intact. |
| P10 | Female | 71 | 20/60 | 20/100 | I have Glaucoma. My field of vision is 5% eyesight. 5% in my left and 5% in the right remaining. |
5.2. Apparatus
The study was conducted in a well-lit lab environment. Participants sat in front of a large table, where we placed nine different kitchen tools—knife, spoon, fork, scissors, ladle, spatula, pan/pot, cup, and carafe. We used a dark green table cloth for the experiment table to produce low contrast. We also prepared a yellow wooden cutting board, a bowl, a piece of cheese, and a stick of butter for the participants to use in the study, however, CookAR can only recognize and augment the nine aforementioned kitchen tools at the current stage. Lastly, we recorded the experiment using a separate laptop and a smartphone on a tripod.
5.3. Procedure
The single-session 90-minute study consisted of three phases. In Part 1, we asked participants to grab cooking utensils in three conditions: (1) without CookAR, (2) with CookAR displaying augmentations for whole objects, and (3) with CookAR displaying affordance augmentations (Figure 5). In Part 2, participants completed a full cooking task where they made macaroni and cheese while using CookAR. Finally, in Part 3, participants brainstormed tool affordances and desired augmentation designs while examining and critiquing design probes. Prior to the study tasks, participants completed a tutorial to become more familiar with CookAR. They then interacted with our AR device for 30 minutes: a 5-minute tutorial, 10 minutes of object grabbing, and 15 minutes of free-form cooking. We provide more details below.
Tutorial. Participants were asked to first complete a tutorial task: interact with a cooking pot using CookAR. Participants wore and adjusted the Oculus Quest 2 headset and became acquainted with CookAR and its affordance augmentations. Once participants achieved a comfortable fit and an understanding of CookAR, the study proceeded.
Part 1: Comparison Study. In Part 1, participants were asked to locate and pick up cooking tools under three conditions: (1) without CookAR, (2) with CookAR displaying whole object augmentations, and (3) with CookAR displaying affordance augmentations (See Figure 5). We counterbalanced the condition order via Latin Square. In each condition, participants conducted five trials of tool-picking-up tasks. We randomly chose a cooking tool per trial from the nine kitchen tools on the experiment table (See Section 5.2). Participants were asked to keep their eyes closed until the researcher named an object to reduce the effect of memory on task performance. Moreover, the researcher rearranged the placement and angle of the cooking tools between each condition. After each condition, we asked participants three 7-point Likert questions about effectiveness, comfort, and distraction, as well as open-ended questions regarding their experience with each augmentation condition. After all 15 trials, we asked participants to compare the pros and cons of the three conditions, suggest improvements for CookAR, and identify potential applications of CookAR outside of kitchen contexts.
Part 2: Full Cooking Task. In Part 2 of the study, we asked participants to cook a macaroni and cheese dish using ARCook with affordance augmentations. We provided participants with step-by-step instructions to ensure that participants interact with all nine objects ARCook can recognize: (1) grab a cup of water and pour it into a carafe, (2) pour water into a pot using a carafe, (3) cut a piece of butter using a knife, (4) cut a piece of cheese using a pair of scissors, (5) put the macaroni, butter, and cheese into a pot, (6) stir with a spoon, (7) stir with a spatula, (8) place the finished macaroni and cheese in a bowl using a ladle, and (9) pick up a fork and enjoy. For the safety of our participants, we supplied a knife with a dull edge and avoided the use of heat. As participants completed this task, they were encouraged to think aloud, articulating how the affordance augmentations support or hinder their activities, how CookAR impacted their overall cooking experience, and any suggestions they had for augmentation designs. After completing the free-form cooking task, we asked participants to reflect on these same topics through open-ended questions.
Part 3: Brainstorming and Co-Design. In Part 3 of the study, we asked participants to brainstorm future designs and applications of CookAR. We first presented design probes (See Figure 6) and asked participants to critique them. The design probes illustrate various designs, including outlines which reduce visual clutter in comparison to solid-colored overlays, solely displaying either the grabbable or the hazardous augmentation, highlighting the more specific hazardous part such as the sharp edge of a knife blade rather than the whole blade, employing arrows to widen the area covered by the augmentations, and introducing a visual warning system when the user’s hand gets too close to a risky area. The design probes are inspired by prior work on low vision augmentations (Zhao et al., 2016, 2019; Fox et al., 2023; Lang and Machulla, 2021). Subsequently, we invited participants to propose augmentation design ideas for both simple objects like knifes and more intricate objects with more interactive parts beyond grabbable and hazardous areas, like a carafe with many holes and buttons. Lastly, we asked participants to identify scenarios other than a kitchen where CookAR might be beneficial.
5.4. Analysis
We recorded participants’ quotes using Zoom. Transcriptions were first done by the video conferencing software, then the research team manually revised the transcripts. We collected 346 distinct quotes across our 10 LV participants, which we analyzed using reflexive thematic analysis (Braun and Clarke, 2006, 2019). The first author, who facilitated every user study session, created an initial codebook by reviewing the revised study transcripts. The research team then collaboratively iterated on the codebook while checking for bias and coverage. With a final codebook consisting of 23 codes, the first author coded participants’ quotes, after which the team discussed the resulting themes. For likert-scale score analysis, we used a Wilcoxon signed-rank test since the data does not follow a normal distribution.
6. Results
In our three-part qualitative study, participants completed tool-grasping tasks, a free-form cooking task, and a brainstorming session with design probes. Overall, participants found the real-time affordance augmentations helpful when interacting with various cooking tools. They also suggested desired augmentation designs and key affordance parts. We expand on these findings below.
6.1. Affordance vs. Whole Object Augmentations
All but one participant (P6) preferred affordance augmentations over whole object augmentations for supporting kitchen tool interactions. They noted a trade-off between the augmentations’ utility and distraction, with the former generally outweighing the latter: “Seeing one color was less distracting than seeing two colors. But you’d have to know which end of the tool is the handle and which is the working end” (P2). We report participants’ feedback on the effectiveness and distraction of the augmentations below.
Effectiveness. In examining Likert data on perceived effectiveness, the findings showed no significant difference (; ). However, participants on average gave higher ratings to affordance augmentations (5.3; 1.6) over whole object augmentations (4.6; 1.4). Affordance augmentations are advantageous in quickly understanding the overall scene (P1, P10), along with the placement and orientation of individual objects: “It helps to have two colors. I could see that more readily and quickly to understand how to use the object and how it is placed… your system is helpful because where the tool starts and ends and where the handle starts and ends is more clear” (P5). In addition, affordance augmentations become particularly useful when handling objects that have hazardous (9 out of 10 participants) (e.g., sharp, hot ) or small (8/10) (e.g., door handles, buttons on appliances) parts, or have insufficient color contrast (7/10) (e.g., all silver or black cooking tools): “You’ve gotta show parts you can and shouldn’t grab. Green tells me that’s a safe place to go with my hand. Anything not green, I shouldn’t grab… It can help me avoid dangerous parts or perhaps even find small things like remote controllers” (P4). Furthermore, four participants expressed that for some objects with more complex interaction components than “just grab and don’t grab” (P7), like a carafe with its handle, base, buttons, lid, and spout, they would accept the use of more than two colors, although “more than four colors can be quite distracting” (P7). We discuss augmentation designs in Section 6.3.
Distraction & Comfort. While participants qualitatively expressed that the whole object augmentations are less distracting than the affordance augmentations (6/10), the comfort and distraction Likert data was not significantly different (; and ; respectively). The difference in average rating was also negligeable, although participants on average found whole object augmentations to be slightly more comfortable (5.1; 1.3 vs. 5.0; 1.2) and less distracting (2.3; 0.9 vs. 2.5; 1.1). P6, who preferred whole object augmentations during the study, said “I think the more colors you have, the more distracting it becomes. So I prefer just the whole object in green than having 2 or 3 different colors. An outline would be better. I would definitely stay away from multicolored and just stick with one color. I can figure out its different parts.” Additionally, three participants shared that the whole object augmentations could be more useful depending on the scenarios. For tasks such as locating or avoiding objects, where interaction is not the goal, whole object augmentations are more preferable, since they are less distracting: “If I am looking for the remote controller, if it could make the remote stand more out in green or something. I don’t need its parts” (P3).
In summary, if a person’s intent is to interact with an object, affordance augmentations are more helpful than whole object augmentations. Conversely, in cases where interaction is not the objective, whole object augmentations may be preferred as they are less distracting. In our study, participants had to interact with various kitchen tools, hence most of them found our affordance augmentations more useful.
6.2. Experiences with CookAR
All participants were able to complete all free-form cooking steps within three to five minutes. However, due to current technical limitations in accurately segmenting object affordances, deploying heavy CV models, and rendering spatially accurate overlays in real-time on AR headsets, CookAR experienced recognition errors and perceived latency, which affected LV users’ experiences. All participants observed “flickering” and inaccurate affordance augmentations. Participants also pointed out that “the colors took some time to catch up” (P3) as they quickly rotated their head. Nonetheless, all but P6 saw potential in CookAR to assist with kitchen tool interactions and beyond: “I like the contrasting color. I just wish it more closely matched the object’s actual location. I think this highlighting scene is a great start. If the system is perfect, the dual color highlight system would be great and most useful. The system would be perfect if I am in a kitchen or just trying to grab really anything” (P1). P1, P5, P9, and P10 were particularly excited as they were able to better visually perceive object information: “This is fun! I can also use my eyes more to see shape and how it can be used. I want to try your system again in the future once you make it better” (P9).
Recognizing the potential of CookAR, participants identified additional use cases of a CookAR-like system that go beyond a cooking scenario, including cleaning (P3, P6, P7, P9), woodworking (P4, P5, P9), walking outdoors (P2, P6, P9), driving (P4, P6, P7), visiting a foreign kitchen (P2, P5), restaurants (P3, P9), gardening (P5, P10), watching sports (P7, P10), playing board games (P7, P10), going down stairs (P7, P9), identifying pill bottles (P7), and interacting with appliances with multiple buttons like a toaster (P1).
6.3. Desired Augmentation Designs
We report participants’ preferences on augmentation designs for grabbable and hazardous areas based on the design probes.
Combining solid and outline augmentations. As opposed to solid-colored overlays, nine participants preferred a mix of solid and outline augmentations, because solid colors are more salient, whereas outlines are less distracting: “Solid colors are helpful because they grab my attention… outlines are helpful because I can still see the part I’m trying to use with less distraction” (P5). Among those nine participants, all but P7 preferred solid-colored overlays for the grabbable area and outlines for the hazardous area because “the grabbable area is the most important” (P3, P4, P5, P8, P9), “all you need to know is its shape” (P4, P9), and “other parts should be outlined since you may want to do more with it, and solid color just makes it harder to use it” (P4, P8, P9). However, P7 preferred the outline for the grabbable area since it is not distracting and still shows the shape of the handle.
For the risky area to outline, P8 preferred highlighting solely the exact hazard (e.g., the sharp edge of the knife blade), as opposed to the entire dangerous part of an object (e.g., the whole knife blade), because she needs to know the relatively safer area for interaction. For example, as she described, “I might grab the top of the blade when I’m dicing or chopping. This tells you exactly where you shouldn’t touch.” In contrast, all other participants wanted the outline augmentation because it is less distracting, yet still defines the overall shape of the hazardous area (P4, P5, P7, P9, P10) and what it is used for (P4, P7, P9). For example, P9 said, “I prefer to see the outlines on the [whole] blade, just so that way, you know which type of blade you’re grabbing. Cause a bread knife would look different from your knife. Some are thinner, some are fatter. People can be quite picky about their knife choices.” P3, who preferred to outline dangerous parts, cautioned that “depending on how low vision you are on the spectrum, you may need solid colors to be able to see some of these objects.”
Lastly, P1, P4, and P10 expressed concerns that overlaying perfectly aligned solid-colored affordance augmentations can be technically challenging. They suggested a colored circle may be enough, since they only need “a hint to see a glimmer of the object” and determine how the objects are oriented (P1).
Enhancing color contrast. Using colors to distinguish object affordances was well-received, as participants often color code their own cooking tools: “So I always try to get things color coded… especially if things are in drawers, it takes a lot of cognition for me to tell you what’s what. If it’s colored, it’s so much easier. This system is huge cause it’s doing color coding for me” (P1).
Every participant favored using green for safe-to-grab and red for dangerous areas, as “green signals ‘yes’ while ‘red’ signals no” (P4). However, P4, P7, and P9 struggled to clearly see our choice of red and requested a brighter shade of red, with P4 even suggesting white. Moreover, participants noted that the color contrast between the tool and the background is more important than the specific colors used, since a lot of objects in a kitchen are white, silver, or black with low contrast. For example, when cooking mac and cheese in the study, most (8/10) participants found it challenging to cut butter and understand where the yellow butter starts and ends because it was on a yellow wooden cutting board. To address this, P3, P4, P5, and P10 suggested the system should automatically select colors that contrast against the background: “The background you have it against will make a big difference, right? So on a darker background, I should be getting light colors” (P4). P7 jokingly said, “I mean, a green stick of butter could be weird, but it would at least let me cut better.”
Auditory feedback. Instead of visual augmentations, all participants preferred auditory feedback for warning in urgent scenarios (e.g., when the user’s hands get too close to a knife blade), as a visual warning could be easily missed by low vision users (P3, P5, P7, P9) and also makes the overall visual field busier (P4, P10). P3, P5, and P6 suggested small yet noticeable audio such as “beep beep,” while P7 and P10 preferred explicit verbal waning (e.g., “stop”) since small noises can also be generated by other devices, such as a microwave or a fridge. P4 and P9 further suggested the system should employ different auditory signals for different hazards.
Action-aware Augmentations. As opposed to constantly augmenting all affordances, half of the participants suggested generating augmentations based on users’ current tasks or behaviors to reduce potential distraction. For example, with a knife, both the handle and blade can be augmented to start, then when a person grabs it, the handle augmentations could be turned off (P7); or, as a person gets close to a carafe with a cup of water, the rim of the carafe could be highlighted (P4). Moreover, seven out of 10 participants also suggested using voice commands to control the augmentations, such as turning on and off an augmentation or adjusting the augmentation design (e.g., colors or forms).
6.4. Additional Tool Affordances
In addition to the grabbable and hazardous affordances focused in our CookAR system, participants collectively suggested five other important affordances for kitchen tools: (1) entry area, (2) exit area, (3) containment area, (4) intersection area, and (5) activation area. We elaborate on these seven affordances along with participants’ preferred augmentation designs.
Grabbable area. A grabbable area is the part designed for safe handling or manipulation. This can include handles, grips, or any part intended for direct hand contact. For grabbable part of an object, participants preferred green solid-colored augmentations (Section 6.3).
Hazardous area. A hazard area is the part that poses potential risks or dangers to the user. This could include sharp edges, hot surfaces, or any part that can cause injury if touched or mishandled. For hazardous part of an object, participants preferred red outline augmentations (Section 6.3).
Entry area. An entry area is the part designed for initiating access, such as pouring. This could be the rim of a cup or a pot, the opening of a carafe, or any designated point that allows entry into an object’s containment space. All participants consistently noted that this area on an object should be augmented by an outline rather than a solid color, as the latter obstructs relevant actions like pouring or scooping: “The color blobs hide the item that you’re trying to put things into virtually completely. And so I can’t really tell if I am pouring something in correctly” (P4).
Exit area. An exit area is defined as the point through which contents are meant to be released. This could be the spouts, holes, or any defined pathway that guides content out of the object’s containment space, and it can be the same as the entry area for some objects, such as bowls and cups. Several participants (4/10) suggested that the carafe’s spout, similar to the entry area, should be outlined: “Highlighting the spout would be helpful if you had to pour, because if I poured in the wrong place, I wouldn’t know until something spills. I think I can pour more effectively if you highlighted this by aligning it with the edge of a cup or something. An outline would be great so I can see the water flowing out” (P7).
Containment area. A containment area has some depth and is meant to hold content within, such as food and liquid. This could be the interior of a cup or pot, the base of a spoon or ladle, or any defined space within the object that is meant to keep something in. The current solid-colored overlays in CookAR interfere with visibility of the containment space. Instead, all participants wanted CookAR to augment only the entry and exit areas using outlines, leaving the containment space without any augmentation.
Additionally, eight participants expressed that they need assistance with understanding the depth of the containment space and the amount of content it already holds. As P8 expressed, “A lot of people with low vision cannot see inside and know how much water they can pour. So somehow showing the water level and size of the teapot [is helpful]. Mine is a lot bigger, it makes 12 cups or something, and it’s all black, so it’s even harder to see what’s inside.” While P9 has a strategy to overcome this challenge by using her finger to feel the liquid level, she cannot use it when the water is hot. She thus suggested the system rendering “a blue disk” to indicate the water level. In terms of augmenting the depth of the containment area, participants suggested using a virtual line from the rim to the bottom of the pot (P3, P4, P5, P9, P10), a measuring tape with ticks (P4, P5, P9, P10), or a line with changing colors (e.g., a green line that turns red as water fills up), (P4, P5). P10 further suggested an auditory cue (e.g., a ‘ding’ sound) to indicate action milestones, such as when reached quarter of a cup.
Intersection area. An intersection area is where parts of two or more objects meet. This could be where a knife blade touches the butter for cutting or where a cup touches a pot for pouring. Interactions that require precise alignment between two objects are particularly challenging to our LV participants. Half of the participants suggested generating augmentations to highlight the intersections or relationships between two interacting objects, for example, the location where a knife cuts the butter (P5) or the alignment between a ladle and a bowl when pouring (P9). As P9 mentioned, “Using a ladle has always been a problem for me. Pouring the ladle into things is usually the hardest part, because you never know if the ladle is in the right spot or too wide out of the way. Maybe, if you have the ladle on top of a bowl, [CookAR should render] a [virtual] shadow that gets casted onto the bowl.”
Activation area. An activation area is designed for initiating, activating, or turning on an object’s function or features. This could be buttons, switches, touch-sensitive surfaces, or any interactive components that trigger the operation of an object. Participants identified activation areas on many household appliances, such as buttons and dials on stove tops, microwaves, or coffee pots. They are used for various purposes including starting a machine, opening a lid, and adjusting settings. Seven participants preferred outline augmentations for the activation area: “I just bought a vacuum with multiple buttons. You would want different colors for the handles and buttons” (P8). P10 further suggested a clock-like augmentation in addition to the outline for turnable dials: “On a stove, I don’t know what is medium heat. As I turn the knobs on a stove, the system could show me ‘2 o’clock,’ ‘3 o’clock,’ and so on. ‘6 o’ clock’ is probably a medium heat. ‘9 o’clock’ is probably a high heat.”
7. Discussion
CookAR explores the use of affordance augmentations to enhance kitchen tool interactions for LV people and advances the state-of-the-art in AI-powered AR systems. Results from our user study indicates a preference for affordance augmentations over whole object during tool interactions. Additionally, participants favored augmentation designs that incorporate both solid-colored and outlined overlays with contrasting colors. We discuss design implications for affordance augmentations as well as current limitations and future opportunities of AI-powered AR systems for low vision.
7.1. Design Implications for Affordances
Throughout the study, our LV participants proposed a range of affordances for kitchen tools and indicated their preferred augmentation designs for each. In this section, we summarize and expand on these suggestions.
When to use affordance augmentations? Our study indicates that visual augmentations should maximize utility and minimize visual clutter and confusion. As such, it is critical to render augmentations tailored to users’ intent and reduce distraction. For example, affordance augmentations that involve multiple pieces and colors would be more preferred to support interactions, while basic whole-object augmentations are more suitable in general visual perception tasks such as avoiding obstacles. Beyond a kitchen, affordance augmentations could also be applied to other scenarios, as our participants suggested, such as gardening, playing board games, and interacting with appliances (reaffirming (Lang and Machulla, 2021)).
Where to apply affordance augmentations? Affordances can refer to any object parts that indicate diverse action or interaction opportunities. However, LV people face distinct interaction challenges, resulting in unique affordances to augment for them. In our qualitative study, we identified seven essential affordances of kitchen tools that represent important yet challenging interaction tasks for LV users. They include: (1) grabbable area, affording touching and handling action; (2) hazardous area, affording risks and avoidance; (3) entry, affording a target to aim at or pour in; (4) exit, affording pouring out and usually requiring accurate alignment with the entry of another object (e.g., food transferring or pouring); (5) containment area, affording holding content in and preferring augmentations on the content amount (e.g., ingredient measurement); (6) intersection area, affording touching or interaction between two objects; and (7) activation area, affording control features on an object. This affordance framework summarizes the critical areas on objects as well as the hand-object (e.g., grabbable vs. hazardous areas) and object-object (e.g., entry-exit alignment, intersection between objects) relationships during interactions.
How to augment affordances? Different augmentations should be designed for different types of affordances according to the interaction tasks they indicate. Our study revealed the most preferred augmentations for different affordances. Specifically, solid-colored overlays with great visual emphasis are preferred for grabbable area to enable fast perception and action, measuring augmentations (e.g., line with ticks) for containment area to indicate content amount, and outlines for other affordances to avoid distraction and occlusion. In terms of colors, augmentations should adopt colors with high contrast against the environment. We also suggest leveraging the commonly-agreed meanings of colors, such as green for safe-to-grab areas and red for risky areas. However, while preferred by the LV participants in our study, the green-red combination should be used cautiously considering that LV users may experience color deficiency.
How to control affordance augmentations? Due to the diverse visual abilities and preferences of LV users (Zhao et al., 2015), future systems should support extensive personalization capabilities such as voice-based control for customization and automatic adaptations. For instance, users should be able to adjust different aspects of an augmentation’s design, such as switching it on or off, choosing between solid-colored and outlined overlays, changing the outline’s thickness, and selecting suitable colors. Additionally, these systems could intelligently adapt by recognizing user actions or tasks to only highlight necessary object parts and signal warnings. They can also automatically alter the augmentations’ colors to generate high contrast against the background.
7.2. Challenges in AI-powered AR Development
This paper presents several key technical contributions across CV and HCI by constructing the first egocentric kitchen tool affordance dataset, fine-tuning an affordance segmentation model on our dataset, and developing a fully-functional stereo AR system that generates real-time affordance augmentations. Below, we reflect on key technical challenges stemming from both fields.
AI models for real-world use. Although our fine-tuned model outperforms the base model in accurately segmenting affordances of cooking tools, its mAP is still too low to successfully support dynamic activities like cooking in a real-world deployment study. For instance, its AP@75 is 48.6%, meaning in the worst case, about half of all predictions fail to achieve greater than 75% overlap with the ground truth affordance masks. The recognition results could become even worse in real-world use on a wearable AR system due to the users’ constant head motions and unique behaviors. For example, LV users tend to get much closes to view objects than sighted users (Reynolds et al., 2024). This issue highlights the evaluation gap between HCI and AI: a model that demonstrates a reasonable performance under AI evaluation measures (e.g., mAP, AP@50, AP@75) may be appropriate for an in-lab user study but not a more naturalistic environment found in deployment studies. We suggest that when developing AI models, researchers should consider the potential real-world use cases, human needs, and integration to different hardware platforms (e.g., wearable AR) to enable use in practice.
Affordance models and datasets. As opposed to object recognition models and datasets that attract significant attention in AI (Lin et al., 2014; Deng et al., 2009; He et al., 2016; Liu et al., 2021), the research on affordance models and datasets remains nascent. To address this issue, we collected and labeled an affordance image dataset for kitchen tools and fine-tuned an object detection model on the affordance dataset to balance accuracy and speed. However, due to the relatively small scale of the dataset and the model not being designed for affordance, our system encountered affordance-specific issues. For example, our model often struggled to distinguish different handles, as many handles across various cooking tools look similar. While not affecting the mask generation (so that users can still see the correct augmentations), it interfered with the object tracking model supported by the ZED Mini API, leading to flickering and unstable augmentation rendering. To enhance the system’s robustness and affordance recognition capabilities, future AI research should consider the following key areas: (1) developing larger and more diverse affordance datasets. These datasets should capture a wider variety of object interactions and functionalities, allowing the model to learn from a richer set of scenarios; (2) designing models for affordance detection, such as incorporating training objectives that refine object part relationship understanding for better affordance prediction; (3) on the system level, prioritizing improvements to object tracking algorithms to ensure more stable augmentation rendering, especially in dynamic environments where precise object localization is crucial.
System latency. Latency is always a concern for AR systems, especially considering that off-the-shelf AR devices usually do not have sufficient computational power (e.g., GPU) to support real-time CV models. To enable affordance segmentation, our system streams video data between the AR headset and an external server. However, the latency caused by the data transition prevented the overlays from keeping pace with the participants’ head motions, negatively impacting participants’ trust in CookAR’s intelligence and their perceptions of the system’s usability. To address the system latency, efforts are needed in both software development (e.g., real-time AI models) and hardware advancement (e.g., AR devices with powerful GPUs) to increase the usability of AR systems in dynamic real-world activities.
7.3. Limitations & Future Directions
There are three primary limitations in this work. First, as an initial prototype, we conducted a qualitative user study with a relatively small number of participants to explore usability and solicit reactions to the idea of affordance augmentations in AR. Future work should conduct larger scale studies with more participants and diverse visual conditions. Second, the aforementioned technical challenges and limitations. Third and finally, the current CookAR system provides only one basic affordance augmentation—solid-colored overlays. Building upon the design insights in our study, future work should incorporate more augmentation options (e.g., outlines) and enable more flexible adjustment (e.g., colors, thickness of the outlines), to provide LV users more personalized experience.
8. Conclusion
In this paper, we introduce CookAR, a wearable AR prototype that can overlay affordance augmentations in real-time to support safe and efficient kitchen tool interactions for people with low vision. To build CookAR, we assembled an egocentric kitchen tool affordance dataset, fine-tuned an RTMDet-Ins-l model on our dataset (i.e., RTMDet-Ins-l-Cook), and leveraged this modified model with a stereo depth camera and an AR headset to generate real-time affordance augmentations in 3D space. In a technical evaluation of RTMDet-Ins-l-Cook, we found that it outperforms the base RTMDet-Ins-l model in accurately segmenting affordance masks of cooking tools. We then evaluated the CookAR prototype in a three-part lab study with LV participants. Findings indicate that participants prefer affordance augmentations over whole object augmentations for kitchen tool interactions. Participants expressed a desire for an augmentation design with a combination of solid-colored and outlined overlays. The grabbable area should be augmented using solid green overlays, while other affordances including hazardous, entry, exit, containment, intersection, and activation areas should be outlined using a contrasting color. Participants also asked for enhanced customization and automatic adapations for affordance augmentation designs. Our work suggests future designs of affordance augmentations to enhance kitchen tool interactions for LV people and advances the state-of-the-art in AI-powered AR experiences as low vision aids.
Acknowledgements.
This work was primarily supported by an NSF Graduate Research Fellowship and NSF CHS #1763199.References
- (1)
- Azenkot and Zhao (2017) Shiri Azenkot and Yuhang Zhao. 2017. Designing smartglasses applications for people with low vision. SIGACCESS Access. Comput. 119 (nov 2017), 19–24. https://doi.org/10.1145/3167902.3167905
- Bahl et al. (2023) Shikhar Bahl, Russell Mendonca, Lili Chen, Unnat Jain, and Deepak Pathak. 2023. Affordances from Human Videos as a Versatile Representation for Robotics. CVPR.
- Bilyk et al. (2009) Marie Claire Bilyk, Jessica M Sontrop, Gwen E Chapman, Susan I Barr, and Linda Mamer. 2009. Food experiences and eating patterns of visually impaired and blind people. Canadian Journal of Dietetic practice and research 70, 1 (2009), 13–18.
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative Research in Psychology 3, 2 (2006), 77–101. https://doi.org/10.1191/1478088706qp063oa arXiv:https://www.tandfonline.com/doi/pdf/10.1191/1478088706qp063oa
- Braun and Clarke (2019) Virginia Braun and Victoria Clarke. 2019. Reflecting on reflexive thematic analysis. Qualitative Research in Sport, Exercise and Health 11, 4 (2019), 589–597. https://doi.org/10.1080/2159676X.2019.1628806 arXiv:https://doi.org/10.1080/2159676X.2019.1628806
- Chen et al. (2019) Kai Chen, Jiaqi Wang, Jiangmiao Pang, Yuhang Cao, Yu Xiong, Xiaoxiao Li, Shuyang Sun, Wansen Feng, Ziwei Liu, Jiarui Xu, Zheng Zhang, Dazhi Cheng, Chenchen Zhu, Tianheng Cheng, Qijie Zhao, Buyu Li, Xin Lu, Rui Zhu, Yue Wu, Jifeng Dai, Jingdong Wang, Jianping Shi, Wanli Ouyang, Chen Change Loy, and Dahua Lin. 2019. MMDetection: Open MMLab Detection Toolbox and Benchmark. arXiv preprint arXiv:1906.07155 (2019).
- Chu et al. (2019b) Fu-Jen Chu, Ruinian Xu, Landan Seguin, and Patricio A. Vela. 2019b. Toward Affordance Detection and Ranking on Novel Objects for Real-World Robotic Manipulation. IEEE Robotics and Automation Letters 4, 4 (2019), 4070–4077. https://doi.org/10.1109/LRA.2019.2930364
- Chu et al. (2019a) Fu-Jen Chu, Ruinian Xu, and Patricio A. Vela. 2019a. Learning Affordance Segmentation for Real-World Robotic Manipulation via Synthetic Images. IEEE Robotics and Automation Letters 4, 2 (2019), 1140–1147. https://doi.org/10.1109/LRA.2019.2894439
- Chuang et al. (2018) Ching-Yao Chuang, Jiaman Li, Antonio Torralba, and Sanja Fidler. 2018. Learning to Act Properly: Predicting and Explaining Affordances from Images. In 2018 IEEE/CVF Conference on Computer Vision and Pattern Recognition. 975–983. https://doi.org/10.1109/CVPR.2018.00108
- Corn and Erin (2010) Anne Lesley Corn and Jane N Erin. 2010. Foundations of low vision: Clinical and functional perspectives. American Foundation for the Blind.
- Damen et al. (2018) Dima Damen, Hazel Doughty, Giovanni Maria Farinella, Sanja Fidler, Antonino Furnari, Evangelos Kazakos, Davide Moltisanti, Jonathan Munro, Toby Perrett, Will Price, and Michael Wray. 2018. Scaling Egocentric Vision: The EPIC-KITCHENS Dataset. In Proceedings of the European Conference on Computer Vision (ECCV).
- Deemer et al. (2018) Ashley D Deemer, Christopher K Bradley, Nicole C Ross, Danielle M Natale, Rath Itthipanichpong, Frank S Werblin, and Robert W Massof. 2018. Low vision enhancement with head-mounted video display systems: are we there yet? Optometry and Vision Science 95, 9 (2018), 694–703.
- Deng et al. (2009) Jia Deng, Wei Dong, Richard Socher, Li-Jia Li, Kai Li, and Li Fei-Fei. 2009. Imagenet: A large-scale hierarchical image database. In 2009 IEEE conference on computer vision and pattern recognition. Ieee, 248–255.
- Everingham et al. (1999) MR Everingham, BT Thomas, T Troscianko, et al. 1999. Head-mounted mobility aid for low vision using scene classification techniques. The International Journal of Virtual Reality 3, 4 (1999), 3.
- Findlater et al. (2019) Leah Findlater, Bonnie Chinh, Dhruv Jain, Jon Froehlich, Raja Kushalnagar, and Angela Carey Lin. 2019. Deaf and Hard-of-hearing Individuals’ Preferences for Wearable and Mobile Sound Awareness Technologies. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (Glasgow, Scotland Uk) (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3290605.3300276
- Fox et al. (2023) Dylan R. Fox, Ahmad Ahmadzada, Clara Tenia Wang, Shiri Azenkot, Marlena A. Chu, Roberto Manduchi, and Emily A. Cooper. 2023. Using augmented reality to cue obstacles for people with low vision. Opt. Express 31, 4 (Feb 2023), 6827–6848. https://doi.org/10.1364/OE.479258
- Garrido-Vásquez and Schubö (2014) Patricia Garrido-Vásquez and Anna Schubö. 2014. Modulation of visual attention by object affordance. Frontiers in Psychology 5 (2014), 70664.
- Gibson (1966) James Jerome Gibson. 1966. The senses considered as perceptual systems. (1966).
- Gibson (1977) James J Gibson. 1977. The theory of affordances. Hilldale, USA 1, 2 (1977), 67–82.
- Gibson (2014) James J Gibson. 2014. The ecological approach to visual perception: classic edition. Psychology press.
- Gupta et al. (2018) Anshul Gupta, Juraj Mesik, Stephen A Engel, Rebecca Smith, Mark Schatza, Aurélie Calabrese, Frederik J Van Kuijk, Arthur G Erdman, and Gordon E Legge. 2018. Beneficial effects of spatial remapping for reading with simulated central field loss. Investigative ophthalmology & visual science 59, 2 (2018), 1105–1112.
- He et al. (2017) Kaiming He, Georgia Gkioxari, Piotr Dollár, and Ross Girshick. 2017. Mask r-cnn. In Proceedings of the IEEE international conference on computer vision. 2961–2969.
- He et al. (2016) Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun. 2016. Deep residual learning for image recognition. In Proceedings of the IEEE conference on computer vision and pattern recognition. 770–778.
- Hicks et al. (2013) Stephen L Hicks, Iain Wilson, Louwai Muhammed, John Worsfold, Susan M Downes, and Christopher Kennard. 2013. A depth-based head-mounted visual display to aid navigation in partially sighted individuals. PloS one 8, 7 (2013), e67695.
- Huang et al. (2019) Jonathan Huang, Max Kinateder, Matt J Dunn, Wojciech Jarosz, Xing-Dong Yang, and Emily A Cooper. 2019. An augmented reality sign-reading assistant for users with reduced vision. PloS one 14, 1 (2019), e0210630.
- Hwang and Peli (2014) Alex D Hwang and Eli Peli. 2014. An augmented-reality edge enhancement application for Google Glass. Optometry and vision science 91, 8 (2014), 1021–1030.
- Jain et al. (2018a) Dhruv Jain, Bonnie Chinh, Leah Findlater, Raja Kushalnagar, and Jon Froehlich. 2018a. Exploring Augmented Reality Approaches to Real-Time Captioning: A Preliminary Autoethnographic Study. In Proceedings of the 2018 ACM Conference Companion Publication on Designing Interactive Systems (Hong Kong, China) (DIS ’18 Companion). Association for Computing Machinery, New York, NY, USA, 7–11. https://doi.org/10.1145/3197391.3205404
- Jain et al. (2015) Dhruv Jain, Leah Findlater, Jamie Gilkeson, Benjamin Holland, Ramani Duraiswami, Dmitry Zotkin, Christian Vogler, and Jon E. Froehlich. 2015. Head-Mounted Display Visualizations to Support Sound Awareness for the Deaf and Hard of Hearing. In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems (Seoul, Republic of Korea) (CHI ’15). Association for Computing Machinery, New York, NY, USA, 241–250. https://doi.org/10.1145/2702123.2702393
- Jain et al. (2018b) Dhruv Jain, Rachel Franz, Leah Findlater, Jackson Cannon, Raja Kushalnagar, and Jon Froehlich. 2018b. Towards Accessible Conversations in a Mobile Context for People who are Deaf and Hard of Hearing. In Proceedings of the 20th International ACM SIGACCESS Conference on Computers and Accessibility (Galway, Ireland) (ASSETS ’18). Association for Computing Machinery, New York, NY, USA, 81–92. https://doi.org/10.1145/3234695.3236362
- Jocher et al. (2023) Glenn Jocher, Ayush Chaurasia, and Jing Qiu. 2023. Ultralytics YOLO. https://github.com/ultralytics/ultralytics
- Jones et al. (2019) Nabila Jones, Hannah Elizabeth Bartlett, and Richard Cooke. 2019. An analysis of the impact of visual impairment on activities of daily living and vision-related quality of life in a visually impaired adult population. British Journal of Visual Impairment 37, 1 (2019), 50–63.
- Kashyap (2020) Avyay Ravi Kashyap. 2020. Behaviors, Problems and strategies of visually impaired persons during meal preparation in the Indian context: challenges and opportunities for Design. In Proceedings of the 22nd International ACM SIGACCESS Conference on Computers and Accessibility. 1–3.
- Kim et al. (2022) Minyung Kim, Sooyoung Hwang, Kyoungmin Choi, Youkeun Oh, and Dokshin Lim. 2022. Vision-Based Cooking Assistance System for Visually Impaired People. In International Conference on Human-Computer Interaction. Springer, 540–547.
- Kirillov et al. (2023) Alexander Kirillov, Eric Mintun, Nikhila Ravi, Hanzi Mao, Chloe Rolland, Laura Gustafson, Tete Xiao, Spencer Whitehead, Alexander C. Berg, Wan-Yen Lo, Piotr Dollár, and Ross Girshick. 2023. Segment Anything. arXiv:2304.02643 (2023).
- Kwon et al. (2012) MiYoung Kwon, Chaithanya Ramachandra, PremNandhini Satgunam, Bartlett W Mel, Eli Peli, and Bosco S Tjan. 2012. Contour enhancement benefits older adults with simulated central field loss. Optometry and vision science 89, 9 (2012), 1374–1384.
- Lang and Machulla (2021) Florian Lang and Tonja Machulla. 2021. Pressing a Button You Cannot See: Evaluating Visual Designs to Assist Persons with Low Vision through Augmented Reality. In Proceedings of the 27th ACM Symposium on Virtual Reality Software and Technology (Osaka, Japan) (VRST ’21). Association for Computing Machinery, New York, NY, USA, Article 39, 10 pages. https://doi.org/10.1145/3489849.3489873
- Lee et al. (2023) Jaewook Lee, Devesh P. Sarda, Eujean Lee, Amy Lee, Jun Wang, Adrian Rodriguez, and Jon E. Froehlich. 2023. Towards Real-time Computer Vision and Augmented Reality to Support Low Vision Sports: A Demonstration of ARTennis. In Adjunct Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (, San Francisco, CA, USA,) (UIST ’23 Adjunct). Association for Computing Machinery, New York, NY, USA, Article 81, 3 pages. https://doi.org/10.1145/3586182.3615815
- Li et al. (2021) Franklin Mingzhe Li, Jamie Dorst, Peter Cederberg, and Patrick Carrington. 2021. Non-Visual Cooking: Exploring Practices and Challenges of Meal Preparation by People with Visual Impairments. In Proceedings of the 23rd International ACM SIGACCESS Conference on Computers and Accessibility (, Virtual Event, USA,) (ASSETS ’21). Association for Computing Machinery, New York, NY, USA, Article 30, 11 pages. https://doi.org/10.1145/3441852.3471215
- Li et al. (2024) Franklin Mingzhe Li, Michael Xieyang Liu, Shaun K Kane, and Patrick Carrington. 2024. A Contextual Inquiry of People with Vision Impairments in Cooking. arXiv preprint arXiv:2402.15108 (2024).
- Lin et al. (2014) Tsung-Yi Lin, Michael Maire, Serge Belongie, James Hays, Pietro Perona, Deva Ramanan, Piotr Dollár, and C Lawrence Zitnick. 2014. Microsoft coco: Common objects in context. In Computer Vision–ECCV 2014: 13th European Conference, Zurich, Switzerland, September 6-12, 2014, Proceedings, Part V 13. Springer, 740–755.
- Liu et al. (2021) Ze Liu, Yutong Lin, Yue Cao, Han Hu, Yixuan Wei, Zheng Zhang, Stephen Lin, and Baining Guo. 2021. Swin transformer: Hierarchical vision transformer using shifted windows. In Proceedings of the IEEE/CVF international conference on computer vision. 10012–10022.
- Loshin and Juday (1989) David S Loshin and Richard D Juday. 1989. The programmable remapper: clinical applications for patients with field defects. Optometry and Vision Science 66, 6 (1989), 389–395.
- Luddecke and Worgotter (2017) Timo Luddecke and Florentin Worgotter. 2017. Learning to segment affordances. In Proceedings of the IEEE International Conference on Computer Vision Workshops. 769–776.
- Luo and Peli (2006) Gang Luo and Eli Peli. 2006. Use of an augmented-vision device for visual search by patients with tunnel vision. Investigative ophthalmology & visual science 47, 9 (2006), 4152–4159.
- Luo et al. (2022) Hongchen Luo, Wei Zhai, Jing Zhang, Yang Cao, and Dacheng Tao. 2022. Learning Affordance Grounding from Exocentric Images. In CVPR.
- Lyu et al. (2022) Chengqi Lyu, Wenwei Zhang, Haian Huang, Yue Zhou, Yudong Wang, Yanyi Liu, Shilong Zhang, and Kai Chen. 2022. RTMDet: An Empirical Study of Designing Real-Time Object Detectors. arXiv:2212.07784 [cs.CV]
- Malu and Findlater (2014) Meethu Malu and Leah Findlater. 2014. ”OK Glass?” A Preliminary Exploration of Google Glass for Persons with Upper Body Motor Impairments. In Proceedings of the 16th International ACM SIGACCESS Conference on Computers & Accessibility (Rochester, New York, USA) (ASSETS ’14). Association for Computing Machinery, New York, NY, USA, 267–268. https://doi.org/10.1145/2661334.2661400
- Malu and Findlater (2015) Meethu Malu and Leah Findlater. 2015. Personalized, Wearable Control of a Head-mounted Display for Users with Upper Body Motor Impairments. In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems (Seoul, Republic of Korea) (CHI ’15). Association for Computing Machinery, New York, NY, USA, 221–230. https://doi.org/10.1145/2702123.2702188
- Massof and Rickman (1992) Robert W Massof and Douglas L Rickman. 1992. Obstacles encountered in the development of the low vision enhancement system. Optometry and vision science 69, 1 (1992), 32–41.
- Massof et al. (1994) Robert W Massof, Douglas L Rickman, Peter A Lalle, et al. 1994. Low vision enhancement system. Johns Hopkins APL Technical Digest 15, 2 (1994), 120–125.
- McNaney et al. (2014) Roisin McNaney, John Vines, Daniel Roggen, Madeline Balaam, Pengfei Zhang, Ivan Poliakov, and Patrick Olivier. 2014. Exploring the acceptability of google glass as an everyday assistive device for people with parkinson’s. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Toronto, Ontario, Canada) (CHI ’14). Association for Computing Machinery, New York, NY, USA, 2551–2554. https://doi.org/10.1145/2556288.2557092
- Miller et al. (2017) Ashley Miller, Joan Malasig, Brenda Castro, Vicki L. Hanson, Hugo Nicolau, and Alessandra Brandão. 2017. The Use of Smart Glasses for Lecture Comprehension by Deaf and Hard of Hearing Students. In Proceedings of the 2017 CHI Conference Extended Abstracts on Human Factors in Computing Systems (, Denver, Colorado, USA,) (CHI EA ’17). Association for Computing Machinery, New York, NY, USA, 1909–1915. https://doi.org/10.1145/3027063.3053117
- Mills et al. (2017) Susanna Mills, Martin White, Heather Brown, Wendy Wrieden, Dominika Kwasnicka, Joel Halligan, Shannon Robalino, and Jean Adams. 2017. Health and social determinants and outcomes of home cooking: A systematic review of observational studies. Appetite 111 (2017), 116–134.
- Miyawaki et al. (2012) Kenzaburo Miyawaki, Mutsuo Sano, Syunichi Yonemura, and Michiko Ode. 2012. Cooking rehabilitation support for self-reliance of cognitive dysfunction patients. In Proceedings of the ACM multimedia 2012 workshop on Multimedia for cooking and eating activities. 19–24.
- Montero (2005) Pilar Montero. 2005. Nutritional assessment and diet quality of visually impaired Spanish children. Annals of Human Biology 32, 4 (2005), 498–512.
- Mynatt and Rogers (2001) Elizabeth D Mynatt and Wendy A Rogers. 2001. Developing technology to support the functional independence of older adults. Ageing International 27, 1 (2001), 24–41.
- Nguyen et al. (2016) Anh Nguyen, Dimitrios Kanoulas, Darwin G. Caldwell, and Nikos G. Tsagarakis. 2016. Detecting object affordances with Convolutional Neural Networks. In 2016 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS). 2765–2770. https://doi.org/10.1109/IROS.2016.7759429
- Nguyen et al. (2017) Anh Nguyen, Dimitrios Kanoulas, Darwin G. Caldwell, and Nikos G. Tsagarakis. 2017. Object-based affordances detection with Convolutional Neural Networks and dense Conditional Random Fields. In 2017 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS). 5908–5915. https://doi.org/10.1109/IROS.2017.8206484
- Olwal et al. (2020) Alex Olwal, Kevin Balke, Dmitrii Votintcev, Thad Starner, Paula Conn, Bonnie Chinh, and Benoit Corda. 2020. Wearable Subtitles: Augmenting Spoken Communication with Lightweight Eyewear for All-day Captioning. In Proceedings of the 33rd Annual ACM Symposium on User Interface Software and Technology (Virtual Event, USA) (UIST ’20). Association for Computing Machinery, New York, NY, USA, 1108–1120. https://doi.org/10.1145/3379337.3415817
- Peli (2001) Eli Peli. 2001. Vision multiplexing: an engineering approach to vision rehabilitation device development. Optometry and Vision Science 78, 5 (2001), 304–315.
- Peng et al. (2018) Yi-Hao Peng, Ming-Wei Hsi, Paul Taele, Ting-Yu Lin, Po-En Lai, Leon Hsu, Tzu-chuan Chen, Te-Yen Wu, Yu-An Chen, Hsien-Hui Tang, and Mike Y. Chen. 2018. SpeechBubbles: Enhancing Captioning Experiences for Deaf and Hard-of-Hearing People in Group Conversations. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (, Montreal QC, Canada,) (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–10. https://doi.org/10.1145/3173574.3173867
- Rector et al. (2015) Kyle Rector, Lauren Milne, Richard E Ladner, Batya Friedman, and Julie A Kientz. 2015. Exploring the opportunities and challenges with exercise technologies for people who are blind or low-vision. In Proceedings of the 17th international ACM SIGACCESS conference on computers & accessibility. 203–214.
- Rehrig et al. (2022) Gwendolyn Rehrig, Madison Barker, Candace E Peacock, Taylor R Hayes, John M Henderson, and Fernanda Ferreira. 2022. Look at what I can do: Object affordances guide visual attention while speakers describe potential actions. Attention, Perception, & Psychophysics 84, 5 (2022), 1583–1610.
- Reynolds et al. (2024) Jarek Reynolds, Chandra Kanth Nagesh, and Danna Gurari. 2024. Salient object detection for images taken by people with vision impairments. In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision. 8522–8531.
- Schipper and Brinkman (2017) Chris Schipper and Bo Brinkman. 2017. Caption Placement on an Augmented Reality Head Worn Device. In Proceedings of the 19th International ACM SIGACCESS Conference on Computers and Accessibility (Baltimore, Maryland, USA) (ASSETS ’17). Association for Computing Machinery, New York, NY, USA, 365–366. https://doi.org/10.1145/3132525.3134786
- Stearns et al. (2017) Lee Stearns, Victor DeSouza, Jessica Yin, Leah Findlater, and Jon E. Froehlich. 2017. Augmented Reality Magnification for Low Vision Users with the Microsoft Hololens and a Finger-Worn Camera. In Proceedings of the 19th International ACM SIGACCESS Conference on Computers and Accessibility (Baltimore, Maryland, USA) (ASSETS ’17). Association for Computing Machinery, New York, NY, USA, 361–362. https://doi.org/10.1145/3132525.3134812
- Stereolabs (2024) Stereolabs. 2024. zed-unity. https://github.com/stereolabs/zed-unity.
- Szpiro et al. (2016b) Sarit Szpiro, Yuhang Zhao, and Shiri Azenkot. 2016b. Finding a store, searching for a product: a study of daily challenges of low vision people. In Proceedings of the 2016 ACM International Joint Conference on Pervasive and Ubiquitous Computing (Heidelberg, Germany) (UbiComp ’16). Association for Computing Machinery, New York, NY, USA, 61–72. https://doi.org/10.1145/2971648.2971723
- Szpiro et al. (2016a) Sarit Felicia Anais Szpiro, Shafeka Hashash, Yuhang Zhao, and Shiri Azenkot. 2016a. How People with Low Vision Access Computing Devices: Understanding Challenges and Opportunities. In Proceedings of the 18th International ACM SIGACCESS Conference on Computers and Accessibility (Reno, Nevada, USA) (ASSETS ’16). Association for Computing Machinery, New York, NY, USA, 171–180. https://doi.org/10.1145/2982142.2982168
- Taylor et al. (2016) Deanna J Taylor, Angharad E Hobby, Alison M Binns, and David P Crabb. 2016. How does age-related macular degeneration affect real-world visual ability and quality of life? A systematic review. BMJ open 6, 12 (2016), e011504.
- Tünnermann et al. (2015) Jan Tünnermann, Norbert Krüger, Bärbel Mertsching, and Wail Mustafa. 2015. Affordance estimation enhances artificial visual attention: Evidence from a change-blindness study. Cognitive computation 7 (2015), 526–538.
- van Rheede et al. (2015) Joram J van Rheede, Iain R Wilson, Rose I Qian, Susan M Downes, Christopher Kennard, and Stephen L Hicks. 2015. Improving mobility performance in low vision with a distance-based representation of the visual scene. Investigative ophthalmology & visual science 56, 8 (2015), 4802–4809.
- Wang et al. (2023a) Ru Wang, Nihan Zhou, Tam Nguyen, Sanbrita Mondal, Bilge Mutlu, and Yuhang Zhao. 2023a. Characterizing Barriers and Technology Needs in the Kitchen for Blind and Low Vision People. arXiv preprint arXiv:2310.05396 (2023).
- Wang et al. (2023b) Ru Wang, Nihan Zhou, Tam Nguyen, Sanbrita Mondal, Bilge Mutlu, and Yuhang Zhao. 2023b. Practices and Barriers of Cooking Training for Blind and Low Vision People. In Proceedings of the 25th International ACM SIGACCESS Conference on Computers and Accessibility (, New York, NY, USA,) (ASSETS ’23). Association for Computing Machinery, New York, NY, USA, Article 57, 5 pages. https://doi.org/10.1145/3597638.3614494
- Washington et al. (2017) Peter Washington, Catalin Voss, Aaron Kline, Nick Haber, Jena Daniels, Azar Fazel, Titas De, Carl Feinstein, Terry Winograd, and Dennis Wall. 2017. SuperpowerGlass: A Wearable Aid for the At-Home Therapy of Children with Autism. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 1, 3, Article 112 (sep 2017), 22 pages. https://doi.org/10.1145/3130977
- Williams et al. (2015) Kristin Williams, Karyn Moffatt, Denise McCall, and Leah Findlater. 2015. Designing Conversation Cues on a Head-Worn Display to Support Persons with Aphasia. In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems (Seoul, Republic of Korea) (CHI ’15). Association for Computing Machinery, New York, NY, USA, 231–240. https://doi.org/10.1145/2702123.2702484
- Zhao et al. (2017) Yuhang Zhao, Michele Hu, Shafeka Hashash, and Shiri Azenkot. 2017. Understanding Low Vision People’s Visual Perception on Commercial Augmented Reality Glasses. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI ’17). Association for Computing Machinery, New York, NY, USA, 4170–4181. https://doi.org/10.1145/3025453.3025949
- Zhao et al. (2019) Yuhang Zhao, Elizabeth Kupferstein, Brenda Veronica Castro, Steven Feiner, and Shiri Azenkot. 2019. Designing AR Visualizations to Facilitate Stair Navigation for People with Low Vision. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology (New Orleans, LA, USA) (UIST ’19). Association for Computing Machinery, New York, NY, USA, 387–402. https://doi.org/10.1145/3332165.3347906
- Zhao et al. (2020) Yuhang Zhao, Elizabeth Kupferstein, Hathaitorn Rojnirun, Leah Findlater, and Shiri Azenkot. 2020. The Effectiveness of Visual and Audio Wayfinding Guidance on Smartglasses for People with Low Vision. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (, Honolulu, HI, USA,) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3313831.3376516
- Zhao et al. (2018) Yuhang Zhao, Elizabeth Kupferstein, Doron Tal, and Shiri Azenkot. 2018. ” It Looks Beautiful but Scary” How Low Vision People Navigate Stairs and Other Surface Level Changes. In Proceedings of the 20th International ACM SIGACCESS Conference on Computers and Accessibility. 307–320.
- Zhao et al. (2015) Yuhang Zhao, Sarit Szpiro, and Shiri Azenkot. 2015. Foresee: A customizable head-mounted vision enhancement system for people with low vision. In Proceedings of the 17th international ACM SIGACCESS conference on computers & accessibility. 239–249.
- Zhao et al. (2016) Yuhang Zhao, Sarit Szpiro, Jonathan Knighten, and Shiri Azenkot. 2016. CueSee: exploring visual cues for people with low vision to facilitate a visual search task. In Proceedings of the 2016 ACM International Joint Conference on Pervasive and Ubiquitous Computing (Heidelberg, Germany) (UbiComp ’16). Association for Computing Machinery, New York, NY, USA, 73–84. https://doi.org/10.1145/2971648.2971730